can you direct me on the right path to solve this problem.
A store sells tents, sleeping bags, an camp stools. A customer buys a tent, 3 sleeping bags and 2 camp stools for $175. The price of the tent is 6 times the cost of a camp stool. The cost of the sleeping bag is $40 more then the cost of a camp stool. I have to find the cost of each item.
Now I know i have to solve for X which will be the price of the stool but I'm unsure what is next.

Answer by stanbon(75887)

A customer buys a tent, 3 sleeping bags and 2 camp stools for $175.
The price of the tent is 6 times the cost of a camp stool.
The cost of the sleeping bag is $40 more then the cost of a camp stool.
----------------
t+3b+2s = 175
t=6s
b=s+40
-----------------
EQUATION:
Substitute to solve for s:
6s+3(s+40)+2s=175
6s+3s+120+2s=175
11s = 55
s = $5 (price of a stool)
t=6s=$30 (price of a tent)
b=s+40=$45 (price of a bag)
==================
Cheers,
Stan H.